The volumes of 3 balloons are in the ratio 5:6:7. By what fractions of themselves must the first two be increased so that the ratio of the volumes may be changed to 7:6:5?
Answer by edjones(7569)

The volumes of 3 balloons are in the ratio 5:6:7.
Let the vol. of the balloons be 5, 6 and 7L respectively
The vol of the last balloon will be unchanged and will be 5/18 of the total volume after the first 2 are inflated more.
5x/18=7 for the 3rd balloon after the 1st 2 have their volumes increased
x=25.2
.
6*25.2/18=8.4L new vol of the 2nd balloon.
.
7*25.2/18=9.8L new vol of the 1st balloon.
.
Fraction of themselves the 1st two must be increased so that the ratio of the volumes may be changed to 7:6:5
8.4/6 = 7/5
9.8/5 = 49/25
.
Ed